OpenStudy (anonymous):

Find the equation of a line that is perpendicular to the line y = 1/3x+8 and contains (-9,0)

jimthompson5910 (jim_thompson5910):

the slope of y = 1/3x+8 is m = ______

OpenStudy (anonymous):

1/3 ?

jimthompson5910 (jim_thompson5910):

flip the fraction to get ____

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

now change the sign to get ____

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

so that's the slop of ANY line perpendicular to y = 1/3x+8

jimthompson5910 (jim_thompson5910):

y = mx+b y = -3x + b ... plug in the perpendicular slope 0 = -3(-9) + b ... plug in the given point (-9,0) now solve for b

jimthompson5910 (jim_thompson5910):

slope*

OpenStudy (anonymous):

b = -27?

jimthompson5910 (jim_thompson5910):

m = -3 b = -27 y = mx+b y = -3x + b y = -3x - 27 That's the equation of the perpendicular line to y = 1/3x+8

OpenStudy (anonymous):

Thank you so much for walking me through it!!

jimthompson5910 (jim_thompson5910):

and that goes through (-9,0)

jimthompson5910 (jim_thompson5910):

you're welcome